OpenStudy (anonymous):

the equation of a circle is (x+2)^2+(y-4)^2=25 whats the equation of the tanget (-5,8)?

OpenStudy (anonymous):

implicit diff for this one \[(x+2)^2+(y-4)^2=25\] \[2(x+2)+2(y-4)y'=0\] \[y'=\frac{x+2}{y-4}\] then replace \(x\) by -5, \(y\) by 8

OpenStudy (anonymous):

i dont get your answer satellite

OpenStudy (anonymous):

you want the slope of the tangent line right?

OpenStudy (anonymous):

yes, but the equation of the slope is 

OpenStudy (anonymous):

that is an equation of the slope if you know two points. you only have one point, the point of tangency so you need the derivative right?

OpenStudy (anonymous):

is this calculus class or something else?

OpenStudy (anonymous):

reason i ask is because i cannot figure out how to find the slope of the tangent line without using calculus

OpenStudy (anonymous):

maybe dpalnc has another method, but i do not

OpenStudy (anonymous):

i did the same too... i thought this was a calc problem.. :) it's an alg2 problem, right?

OpenStudy (anonymous):

find the slope of the line connecting the center to the point of tangency... the slope of your tangent line is the negative reciprocal of that slope.

OpenStudy (anonymous):

well, im in plain geometry, and i tought you find the slope using the equation of the circle

OpenStudy (anonymous):

i did not take calcus yet or algerbra 2 so im stuck

OpenStudy (anonymous):

the equation of the circle is already given. what you need is the center of the circle.

OpenStudy (anonymous):

geometry is good here too...

OpenStudy (anonymous):

yea, that easy, the center of the circle is (-2, 4)

OpenStudy (anonymous):

i dont get how you would find the slope of the tanget from the center of the circle

OpenStudy (anonymous):

good. can you find the slope of the line that connects (-2, 4) with (-5, 8) ?

OpenStudy (anonymous):

okk, its the slope from the center to the tanget is -4/3

OpenStudy (anonymous):

good..

OpenStudy (anonymous):

did you know that a line that is tangent to a circle is ALWAYS perpendicular to the radius at the point of tangency?

OpenStudy (anonymous):

yes haha, but from that how do you find the slope of the tanget?

OpenStudy (anonymous):

how do perpendicular lines relate in terms of their slope?

OpenStudy (anonymous):

i have no idea man that why i need help

OpenStudy (anonymous):

that's the key... perpendicular slopes are negative reciprocals of each other... so if your slope is -4/3, the slope perpendicular to that is.... what?

OpenStudy (anonymous):

ohh! i see, its 3/4!